40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue May 3 21:05:23 2022, yuta, Update, BHD, AS path alignment, removing a lot of green stuff 7x
    Reply  Wed May 4 13:09:42 2022, Anchal, Update, BHD, IMC table balanced again, IMC is locking, YARM is locking 
       Reply  Wed May 4 15:08:59 2022, Tega, Update, BHD, XARM alignment to get flashing IFO_XY_ARMS_Flashing.png
          Reply  Wed May 4 18:52:43 2022, Tega, Update, BHD, POP beam too high at POP SM4 on the ITMX chamber snnp_pr2_pr3_itmy_aligment.png
             Reply  Thu May 5 14:46:22 2022, Anchal, Update, BHD, POP beam height lowered, POP_SM4 raised 
                Reply  Thu May 5 17:05:31 2022, Tega, Update, BHD, IMC & X/Y-arm alignment IFO_aligment_Y_locking_on.png
                Reply  Mon May 9 18:43:03 2022, Anchal, Update, BHD, ITMX table layout corrected 
                   Reply  Mon May 9 18:49:05 2022, Tega, Update, BHD, Relocate green TRX and TRY PDs/cams/optics from PSL table to BS table IMG_20220509_184439943.jpgIMG_20220509_184154722.jpg
                      Reply  Tue May 10 15:46:38 2022, JC, Update, BHD, Relocate green TRX and TRY components from PSL table to BS table IMG_0741.jpegIMG_0753.jpeg
                         Reply  Tue May 10 18:25:43 2022, Paco, Update, BHD, Green Transmission path 
                   Reply  Tue May 10 16:31:14 2022, Tega, Update, BHD, ITMX optlev return beam steered to QPD 
       Reply  Fri May 6 13:48:34 2022, Anchal, Update, BHD, WFS loop instability fixed 
    Reply  Thu May 12 19:55:01 2022, Tega, Update, BHD, AS path alignment AS.JPGOMCchamber.JPG
       Reply  Thu May 12 20:11:18 2022, Anchal, Update, BHD, BHDBS Output beams steered out to ITMY table signal-2022-05-12-201844.jpeg
          Reply  Thu May 12 20:24:29 2022, yuta, Update, BHD, POX and POY investigation POY.JPG
       Reply  Mon May 16 13:22:59 2022, yuta, Update, BHD, REFL and AS paths aligned at AP table REFLPOP.JPGPOPAS.JPG
          Reply  Mon May 16 19:14:17 2022, Anchal, Update, BHD, Camera set on AS path and BHDBS output path 
             Reply  Tue May 17 18:43:38 2022, Anchal, Update, BHD, Placed SRM in ITMY Chamber 
                Reply  Wed May 18 08:30:29 2022, Paco, Update, BHD, SRM OpLev 
                   Reply  Thu May 19 08:51:40 2022, Paco, Update, BHD, SRM OpLev 
                Reply  Wed May 18 09:02:52 2022, Anchal, Update, BHD, WFS1 PD centered 
                Reply  Wed May 18 17:23:15 2022, Anchal, Update, BHD, Placed PRM in BS Chamber 
                   Reply  Thu May 19 18:56:08 2022, yuta, Update, BHD, Current OSEM sensor values with all the suspensions aligned OSEMS.txt
                   Reply  Thu May 19 19:05:59 2022, Anchal, Update, BHD, BS Chamber all work finished, BHD path setup 
                      Reply  Fri May 20 20:03:48 2022, Paco, Update, BHD, ITMY chamber work finished - LO and AS overlapped Screenshot_2022-05-20_20-38-25.pngOSEMs.txt
                         Reply  Mon May 23 13:16:59 2022, Paco, Update, BHD, End of vent - checks OSEMs20220523.txtScreenshot_2022-05-23_13-24-06_IFOaligned_GTRY_still_clipping_but_less.pngPR3position.JPG
                            Reply  Tue May 24 15:21:13 2022, Anchal, Update, BHD, Freeswing tests of new SOS started 
Message ID: 16827     Entry time: Tue May 3 21:05:23 2022     Reply to this: 16829   16848
Author: yuta 
Type: Update 
Category: BHD 
Subject: AS path alignment, removing a lot of green stuff 

[Tega, Yehonathan, Koji, Yuta]

We tried to align AS path this afternoon.
IMC is not aligned now after the work today crying
Green mirrors/perisocope in IMC chamber were removed since some of them was clipping the AS beam, and this changed the balance of the IMC stack and thus MC1 and MC3 alignment.

Summary of changes:
ITMY chamber
  - Rotated AS2 in roll by 90 deg to have more aperture for the transmission (photo)
  - IR beams are now centered on AS1, AS2, AS3 and AS4 (photo, photo)

BS chamber
  - Moved ASL towards -X direction for about 1/4 inch
  - Installed GRY_SM2 at the nominal position (re-used GR_SM3 from IMC chamber)

IMC chamber
  - Removed green optics GR_SM4, GR_SM3, GR_PERI2L (GR_PERI2L is now stored at Xend)
  - Removed IFI camera mirrors FIV1, FIV2 (they are now stored at Xend) (photo, photo)
  - GR_SM4 mount is now reused as GRY_SM1 (Y2-2037-0 is now mounted instead of previously mounted Y2-LW1-2050-UV-45P/AR), and GRY_SM1 is installed at the nominal position (photo)
  - Moved weights to balance the stack

OMC chamber (we don't have OMC in this chamber...)
  - We swapped AS5 and AS6 so that the nobs comes in -X direction to have more spacing between AS beam and IMC REFL beam (photo)
  - Moved weights to balance the stack

What we did:
1. Misaligned ITMX and use ITMY reflected beam to align AS path
2. Centered the IR beam on AS1 using SR2
3. Centered the IR beam on AS2 and ASL using AS1. AS2 was rotated in roll by 90 deg to have more aperture for the transmisson. cool
4. Centered the IR beam on AS3 using AS2 nobs, centered the IR beam on AS4 by rotating AS3 in yaw.
5. "AS beam" (it turned out that what we are looking was actually not the AS beam!! Some stray light) was in +X direction by 1 inch or so at AS5. Moving AS5 to center the beam would clip IMC REFL beam. So we swapped AS5 and AS6 so that the nobs comes in -X direction to have more spacing between AS beam and IMC REFL beam.
6. Balanced OMC chamber stack again using IMC REFL beam as a referece (bring the IMC REFL beam to the reference red circle on the monitor).
7. Tweaked the alignment of TT1 and TT2 to have Yarm flashing to ~0.9 in TRY. angel
8. Moved AS5 towards +X by an inch or so to center the "AS beam."
9. Moved ASL towards -X direction for about 1/4 inch and re-centered the beam by AS1 to see if the "AS beam" gets far from IMC REFL at OMC chamber, but the "AS beam" didn't move much.
10. By blocking the beam from ITMY, we found that "AS beam" was not the actual one. frown
11. Opened IMC chamber and found that AS beam is blocked by the past optics.
12. Removed old green optics and IFI camera mirrors. GR_SM4 mount and GR_SM3 were reused as mentioned above.
13. Tried to balance IMC chamber stack to recover IMC alignment. We used IMC REFL beam as a reference, but it was hard to completely bring the IMC REFL beam to the reference red circle on the monitor. It is now off by a beam diameter or so. No IMC flashing now. crying

Next:
Theoretically, balancing IMC chamber stack would recover all the IFO alignment, but maybe tough. It is maybe easier to align MC1 and MC3 to have IMC locked. Assuming input pointing to IMC is not drifted too much, we should be able to recover Yarm flashing by tweaking TT1 alignment only. However, MC3 SD OSEM is at the edge of the range. We might have to balance the stack more or tweak SD OSEM position.

Attachment 1: log_AS2Rotated.jpg  6.759 MB  Uploaded Tue May 3 23:34:50 2022  | Show | Hide all | Show all
Attachment 2: log_AS5andAS6swapped.JPG  3.925 MB  Uploaded Tue May 3 23:35:20 2022  | Hide | Hide all | Show all
log_AS5andAS6swapped.JPG
Attachment 3: log_IFIMirrorsBeforeRemoving.JPG  3.860 MB  Uploaded Tue May 3 23:35:31 2022  | Hide | Hide all | Show all
log_IFIMirrorsBeforeRemoving.JPG
Attachment 4: log_IFIMirrorsPlacedatXend.JPG  5.035 MB  Uploaded Tue May 3 23:35:51 2022  | Hide | Hide all | Show all
log_IFIMirrorsPlacedatXend.JPG
Attachment 5: log_NewGRY_SM1.JPG  3.697 MB  Uploaded Tue May 3 23:36:05 2022  | Hide | Hide all | Show all
log_NewGRY_SM1.JPG
Attachment 6: log_OnAS3.JPG  3.799 MB  Uploaded Tue May 3 23:36:20 2022  | Hide | Hide all | Show all
log_OnAS3.JPG
Attachment 7: log_OnAS4.JPG  3.635 MB  Uploaded Tue May 3 23:36:32 2022  | Hide | Hide all | Show all
log_OnAS4.JPG
ELOG V3.1.3-